Betty DeRose, Inc. operates two departments, the handling department and
the packaging department. During April, the handling department reported
the following information:

                                           % complete      % complete
                                units         DM           conversion 
work in process, April 1        18,000        38%             71%
units started during April      80,000
work in process, April 30       44,000        28%             74%

The cost of beginning work in process and the costs added during April
were as follows:

                                 DM         Conversion       Total cost
work in process, April 1      $ 64,996       $152,785         $217,781
costs added during April       190,336        232,407          422,743
total costs                    255,332        385,192          640,524

Calculate the total cost of the 18,000 units in beginning work in process
using the FIFO process costing method.
Answer #1
Total cost of the 18,000 units in beginning work in process 269936
Workings:
Direct materials
To complete Work in process, April 1 11160 =18000*(1-38%)
Units started and completed 36000 =80000-44000
Work in process, April 30 12320 =44000*28%
Equivalent units 59480
Conversion
To complete Work in process, April 1 5220 =18000*(1-71%)
Units started and completed 36000 =80000-44000
Work in process, April 30 32560 =44000*74%
Equivalent units 73780
Direct materials Conversion
Costs added during April 190336 232407
Divide by Equivalent units 59480 73780
Cost per Equivalent unit 3.20 3.15
Direct materials Conversion Total
To complete Work in process, April 1 11160 5220
X Cost per Equivalent unit 3.20 3.15
Cost to complete Work in process, April 1 35712 16443 52155
Cost of Work in process, April 1 217781
Cost to complete Work in process, April 1 52155
Total cost of the 18,000 units in beginning work in process 269936